# Text-Guided Multi-Property Molecular Optimization with a Diffusion Language Model

### v1

@@ ABSTRACT.
Molecular optimization (MO) is a crucial stage in drug discovery in which task-oriented generated molecules are optimized to meet practical industrial requirements. Existing mainstream MO approaches primarily utilize external property predictors to guide iterative property optimization. However, learning all molecular samples in the vast chemical space is unrealistic for predictors. As a result, errors and noise are inevitably introduced during property prediction due to the nature of approximation. This leads to discrepancy accumulation, generalization reduction and suboptimal molecular candidates. In this paper, we propose a text-guided multi-property molecular optimization method utilizing transformer-based diffusion language model (TransDLM). TransDLM leverages standardized chemical nomenclature as semantic representations of molecules and implicitly embeds property requirements into textual descriptions, thereby preventing error propagation during diffusion process. Guided by physically and chemically detailed textual descriptions, TransDLM samples and optimizes encoded source molecules, retaining core scaffolds of source molecules and ensuring structural similarities. Moreover, TransDLM enables simultaneous sampling of multiple molecules, making it ideal for scalable, efficient large-scale optimization through distributed computation on web platforms. Furthermore, our approach surpasses state-of-the-art methods in optimizing molecular structural similarity and enhancing chemical properties on the benchmark dataset. The code is available at: https://anonymous.4open.science/r/TransDLM-A901.
Semantic Representation, Multi-Property Molecular Optimization, Diffusion Language Model, Text Guidance
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. INTRODUCTION
Molecule generation has made significant strides with the rapid advancement of generative models, especially conditional generative models  (Hoogeboom et al., 2022; Grisoni et al., 2020). These models have demonstrated promising results during tasks such as drug response prediction (DRP) and drug-target binding affinity (DTA)  (Guan et al., 2023; Huang et al., 2024). Although the molecules generated for specific tasks possess relevant properties, they are still insufficient for application to industrial production. As a result, optimizing generated molecules has become a crucial task, drawing the attention of scientists seeking to improve their usability. However, improving the desired drug candidate properties while retaining the original structural scaffolds is challenging. Moreover, the complexity of enormous chemical space highlights the substantial quantity of molecular candidates. Consequently, traditional molecular optimization (MO) methods primarily depend on the experience, knowledge and intuition of chemists, thus resulting in time-consuming manual labor and reducing the likelihood of finding ideal molecules within a limited time.
@@
To address these challenges, early computational approaches using deep learning generated computational strategies to accelerate the traditional MO paradigm  (Jin et al., 2018; Gupta et al., 2018; Sattarov et al., 2019). These deep learning methods mainly learned from Simplified Molecular Input Line Entry System (SMILES), graphs and three-dimensional (3D) structures  (Weininger, 1988; Mittone et al., 2023; Guo et al., 2021; Yang et al., 2023a). To generate molecules with desired properties, conditional generative models  (Lim et al., 2018; Maziarka et al., 2020; Kotsias et al., 2020) have been adopted as auxiliary controllers for the generative process. However, most of these models concentrate on generating molecules from scratch and optimizing them according to specific rules, thereby neglecting the priority of core scaffold retention during molecular optimization. Consequently, molecules optimized through these methods fail to meet the industrial demands, which requires slight changes to the molecular architecture and substantial increases in the desired properties.
@@
Other widely adopted approaches include guided search-based methods, which aim to find target molecules by exploring compounds’ chemical or latent spaces derived from encoder-decoder models. First, latent space search involves encoding a source molecule into a low-dimensional representation, and exploring its adjacent area to find embeddings that meet the specified constraints  (Jin et al., 2018; Zang and Wang, 2020; Kong et al., 2022). Then, these embeddings are decoded into the chemical space, and a property predictor is used to guide the search process. In contrast, chemical space search methods operate directly within the high-dimensional and discrete chemical space to find molecules that meet given constraints. Various advanced optimization techniques, including reinforcement learning and genetic algorithms, have been applied to this search method.
@@
Despite their proven success, a significant limitation of these guided search-based methods lies in their reliance on external property predictors to iteratively optimize molecular properties. As shown in Fig.  1, external predictors, though effective for estimating target properties, inherently introduce errors and noise due to the nature of approximation. Typically, these methods are trained on finite, often biased datasets and may lack complete generalization in novel chemical spaces, resulting in inaccurate property predictions during the search process. This discrepancy can accumulate over iterations, leading to suboptimal molecular candidates or failure to meet the intended constraints. Moreover, the noise generated by these predictors may cause the search process to deviate from the optimal regions in the chemical or latent spaces, further reducing MO efficiency and effectiveness.
@@
To address the aforementioned MO issues, we explore the utilization of diffusion models in target-oriented MO. Diffusion models  (Ho et al., 2020) have reaped remarkable success in other scientific disciplines, such as computer vision  (Rombach et al., 2022), by generating high-quality data through a gradual denoising process that can capture complex distributions  (Yang et al., 2023b). Accordingly, we propose text-guided multi-property molecular optimization with a transformer-based diffusion language model (TransDLM). This is a novel approach that leverages a diffusion language model to iteratively yield word vectors of molecular SMILES strings and is guided by language descriptions. Due to SMILES strings’ deficiency in explicit semantic clarity of molecular structures and functional groups, we adopt standardized chemical nomenclature as informatively and intuitively molecular semantic representation. To differentiate generation from scratch, we sample molecular word vectors from the token embeddings of source molecules encoded by a pre-trained language model, which collectively and significantly retains the original molecules’ core scaffolds. Instead of relying on an external property predictor, we encode the molecular structure and property information using a pre-trained language model, implicitly embedding property requirements into textual descriptions, guiding diffusion and preventing error propagation. This encoded representation serves as a guiding signal during the diffusion process. By embedding the desired molecular attributes through the language model, the diffusion process is trained on the molecule’s structure and the specific properties we aim to optimize. Additionally, our method is suitable for deployment on web-based platforms due to its high parallel efficiency, enabling large-scale MO across distributed environments. This allows researchers to collaborate and iterate molecular design and optimization processes seamlessly via web interfaces. The TransDLM approach’s advantages are as follows:
@@
  • TransDLM outperforms state-of-the-art methods on the benchmark dataset in optimizing three ADMET  [^1] Absorption, distribution, metabolism, excretion and toxicity properties, LogD, Solubility, and Clearance while maintaining the structural similarity over several metrics.
@@ 2.1. Molecular Optimization
Methods that optimize molecules are classified into three categories: distribution matching-based, guided search-based, and molecular mapping-based approaches.
Distribution Matching-Based Methods.   Previous study  (Xia et al., 2024) has indicated that the matched molecular set (MMS) is used in distribution matching-based methods. In an MMS, the molecular set with poor properties is named the source, and the set with the desired qualities is designated as the target. Molecular distribution matching focuses on learning the mapping relationship between two molecular sets in an MMS dataset. This ensures that the source and target molecular set are adjacent. For instance, Maziarka et al.  (Maziarka et al., 2020) proposed a CycleGAN-based model that encodes each molecule as a point within the latent space by learning the MMS representation distribution. Barshatski et al.  (Barshatski and Radinsky, 2021) designed a multi-property optimization technique called IPCA, which learns a distinct transformation for each optimized property while imposing constraints on the latent embedding space across all transformations. However, distribution matching-based methods are limited, as they fail to optimize specific molecules towards their desired counterparts. Since adjusting the distribution of the whole molecular set is their main goal, these approaches are eventually phased out due to these limitations.
@@
Guided Search-Based Methods.   Guided search-based methods explore target molecules within the molecules’ chemical or latent space learned through encoder-decoder models using molecular property predictors or statistical models for guidance  (Hoffman et al., 2022). For example, Jin et al.  (Jin et al., 2018) introduced a junction tree variational autoencoder (JTVAE), which initially decomposes a molecular graph into a junction tree, with each node representing a molecular substructure. Then, both the junction tree and molecular graph are mapped to latent embeddings. Additionally, a property score predictor is trained in the latent space to cluster molecules with similar properties. Finally, gradient ascent is used to identify embeddings with enhanced property scores within the latent space. Zhou et al.  (Zhou et al., 2019) proposed the MolDQN model, which integrates reinforcement learning with chemical rules to guarantee molecular validity. The model frames molecule modification as a Markov Decision Process (MDP) and uses deep Q-networks  (Hester et al., 2018) to address it, with rewards based on the desired properties and structural similarity constraints. Nevertheless, these guided search-based methods rely on external property predictors to iteratively optimize molecular properties, inevitably introducing errors and noise into the optimization process.
@@
Molecular Mapping-Based Method.   In contrast to MMS, two molecules with moderate structural difference and significant chemical dissimilarities are considered as a matched molecular pair (MMP), where the one with inferior properties is regarded as the source molecule and the other with desired qualities is the target molecule. Molecular mapping involves learning the mapping function $F$ between MMPs. For an MMP $(x,y)$, where $x$ and $y$ exhibit minor differences between consecutive atoms but significant property differences, the molecular mapping is formally denoted as $F(x):x\rightarrow y$. This method is executed using medicinal chemistry transformation rules derived from MMPs, applying the knowledge and expertise of medicinal chemists. For instance, He et al.  (He et al., 2021) designed a model based on Seq2Seq architecture to handle property variations between MMPs. The model uses source molecules and property constraints as inputs and is trained to produce target molecules as outputs. Maragakis et al.  (Maragakis et al., 2020) introduced DESMILES, which utilizes source molecular fingerprints as inputs to generate corresponding SMILES representations. Initially, the transformation from molecular fingerprints to SMILES is pre-trained on a large dataset, and the model is subsequently fine-tuned using the MMPs. Inspired by natural language processing (NLP), Fan et al.  (Fan et al., 2022) developed a semi-supervised approach named BT4MolGen. This method involves training a pseudo-labeled data generation model using MMPs. Then, the generated pseudo-labeled data along with labeled data are employed to train a forward molecular optimization model, addressing the data sparsity issue. Similarly, our TransDLM approach is a molecular mapping-based method that innovatively introduces a diffusion language model into the MO task, resulting in the coherent and effective optimization of molecules.
Figure 2. TransDLM framework. First, our optimization method disassembles the source molecules into fragments using the Recap strategy. Then, it generates textual descriptions utilizing IUPAC names, which guide the denoising process. Meanwhile, the source and target molecules are encoded into embeddings through the SMILES tokenizer and a pre-trained language model, which serve as initial and target vectors, respectively.
@@ 2.2. Language Generation with Diffusion Models
Diffusion models have been extremely successful in generating content in continuous domains, particularly in the images and audio fields  (Yang et al., 2023b). Stable Diffusion  (Rombach et al., 2022) and AudioLDM  (Liu et al., 2023) are categorized as diffusion models in continuous domains. They introduced random noises into latent variables based on latent diffusion models (LDMs) and reverse the process through a series of denoising steps to learn data generation. However, differences in data structure between languages and images exist, since languages are discrete while images exist in continuous domains. In order to tackle this problem, certain approaches retain the discrete nature of text and extend diffusion model to handle discrete data  (Reid et al., 2022; He et al., 2022). On the contrary, other methods utilize embedding layers to map the text to a continuous representation domain, thereby preserving the continuous diffusion steps  (Li et al., 2022; Gong et al., 2022; Yuan et al., 2022; Strudel et al., 2022). Our research aligns with the latter strategy, focusing on word vectors and substantially expanding the functionality and feasibility of diffusion models in SMILES-based MO.
@@ 3.1. Overall Framework
With the solid foundation paved by conditional language generation tasks, exemplified by DiffuSeq  (Gong et al., 2022) and SeqDiffuSeq  (Yuan et al., 2022), exploring text-guided MO has become feasible. In this paper, we aim to optimize a molecule until it matches a specified textual description. Formally, let $C=[w_{0},w_{1},\dots,w_{m}]$ represent the input description, where $w_{i}$ is the $i$-th word in the sequence, and $m$ denotes the text length. Our objective is to develop a model $F$ that takes this text and the original molecule $M_{0}$ as input and produces the corresponding molecule as output, which can be mathematically represented as $M=F(M_{0},C)$. Overall, our TransDLM is composed of four pivotal processes: embedding, noising, denoising, and rounding. As shown in Fig.  2, these processes work in tandem to produce the desired optimized molecules.
The embedding process begins by treating the text sequence $W=[w_{0},w_{1},\dots,w_{n}]$ as an ordered list of words. Each word $w_{i}$ is mapped using an embedding function, yielding $\text{Emb}(W)=[\text{Emb}(w_{0}),\text{Emb}(w_{1}),\dots,\text{Emb}(w_{n})]\in\mathbb{R}^{d\times n}$, where $n$ indicates the sequence length and $d$ represents the embedding dimension. Then, the noising process starts with the matrix $x_{0}$, which is sampled from a Gaussian distribution centered at $\text{Emb}(W)$: $x_{0}\sim\mathcal{N}(\text{Emb}(W),\sigma_{0}\mathbf{I})$.
During the noising process, noise is gradually introduced to $x_{0}$, eventually resulting in pure Gaussian noise $x_{T}\sim\mathcal{N}(0,\mathbf{I})$, where $T$ is a hyperparameter of total diffusion steps. The transition from $x_{t-1}$ to $x_{t}$ is defined as follows:
@@
where $\beta_{t}$ is a hyperparameter representing the amount of noise added at diffusion step $t\in(0,T]$.
Next, the denoising process begins with an encoded source molecule, and sequentially samples $x_{t-1}$ from $x_{t}$, gradually reconstructing the original content. Typically, a neural network is trained to predict $x_{t-1}$ given $x_{t}$. To improve the accuracy of denoising $x_{t}$ towards specific word vectors, a neural network $f_{\theta}$ is trained to predict $x_{0}$ directly from $x_{t}$. Thus, the denoising transition from $x_{t}$ to $x_{t-1}$ can be expressed as:
$p_{\theta}(x_{t-1}\mid x_{t})=$ $\mathcal{N}(x_{t-1};\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}f_{\theta}(x_{t},t)$
(2) $+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}x_{t},\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t}),$
where $\overline{\alpha}_{t}=\prod^{t}_{s=0}(1-\beta_{s})$, facilitating the iterative sampling of $x_{t-1}$ from $x_{t}$ and consequently yielding $x_{0}$.
Finally, the rounding process converts the embedding vectors back into the target molecule’s SMILES string. During this process, each column vector $x_{0}$ is matched to the word with an embedding that has the closest L-2 distance. As a result, the combined denoising and rounding processes effectively transform any initial noise into a coherent SMILES string output.

### v3

@@ ABSTRACT.
Molecular optimization (MO) is a crucial stage in drug discovery in which task-oriented generated molecules are optimized to meet practical industrial requirements. Existing mainstream MO approaches primarily utilize external property predictors to guide iterative property optimization. However, learning all molecular samples in the vast chemical space is unrealistic for predictors. As a result, errors and noise are inevitably introduced during property prediction due to the nature of approximation. This leads to discrepancy accumulation, generalization reduction and suboptimal molecular candidates. In this paper, we propose a text-guided multi-property molecular optimization method utilizing transformer-based diffusion language model (TransDLM). TransDLM leverages standardized chemical nomenclature as semantic representations of molecules and implicitly embeds property requirements into textual descriptions, thereby mitigating error propagation during diffusion process. By fusing physically and chemically detailed textual semantics with specialized molecular representations, TransDLM effectively integrates diverse information sources to guide precise optimization, which enhances the model’s ability to balance structural retention and property enhancement. Additionally, the success of a case study further demonstrates TransDLM’s ability to solve practical problems. Experimentally, our approach surpasses state-of-the-art methods in maintaining molecular structural similarity and enhancing chemical properties on the benchmark dataset.
Drug discovery, Multi-property molecular optimization, Diffusion language model, Text guidance
doi: 10.1016/j.inffus.2025.103907conference: Volume 127, Part C; 1566-2535
@@ 1. Introduction
Molecule generation has made significant strides with the rapid advancement of generative models, especially conditional generative models (Grisoni et al., 2020; Hoogeboom et al., 2022; Li et al., 2025a). These models have demonstrated promising results during tasks such as drug response prediction (DRP) and drug–target binding affinity (DTA) (Guan et al., 2023; Huang et al., 2024; Li et al., 2024). Although the molecules generated for specific tasks possess relevant properties, they are still insufficient for application to industrial production. As a result, optimizing generated molecules has become a crucial task, drawing the attention of scientists seeking to improve their usability. However, improving the desired drug candidate properties while retaining the original structural scaffolds is challenging. Moreover, the complexity of enormous chemical space highlights the substantial quantity of molecular candidates. Consequently, traditional molecular optimization (MO) methods primarily depend on the experience, knowledge and intuition of chemists, thus resulting in time-consuming manual labor and reducing the likelihood of finding ideal molecules within a limited time.
@@
To address these challenges, early computational approaches using deep learning generated computational strategies to accelerate the traditional MO paradigm (Jin et al., 2018; Sattarov et al., 2019). These deep learning methods mainly learned from Simplified Molecular Input Line Entry System (SMILES), graphs and three-dimensional (3D) structures (Weininger, 1988; Wang et al., 2024a; Yu et al., 2025; Chen et al., 2025). To generate molecules with desired properties, conditional generative models (Lim et al., 2018; Maziarka et al., 2020; Kotsias et al., 2020) have been adopted as auxiliary controllers for the generative process. However, most of these models concentrate on generating molecules from scratch and optimizing them according to specific rules, thereby neglecting the priority of core scaffold retention during molecular optimization. Consequently, molecules optimized through these methods fail to meet the industrial demands, which requires slight changes to the molecular architecture and substantial increases in the desired properties.
@@
Other widely adopted approaches include guided search-based methods, which aim to find target molecules by exploring compounds’ chemical or latent spaces derived from encoder–decoder models. First, latent space search involves encoding a source molecule into a low-dimensional representation, and exploring its adjacent area to find embeddings that meet the specified constraints (Jin et al., 2018; Zang and Wang, 2020; Kong et al., 2022). Then, these embeddings are decoded into the chemical space, and a property predictor is used to guide the search process. In contrast, chemical space search methods operate directly within the high-dimensional and discrete chemical space to find molecules that meet given constraints. Various advanced optimization techniques, including reinforcement learning and genetic algorithms, have been applied to this search method.
@@
Despite their proven success, a significant limitation of these guided search-based methods lies in their reliance on external property predictors to iteratively optimize molecular properties. As shown in Fig. 1, external predictors, though effective for estimating target properties, inherently introduce errors and noise due to the nature of approximation. Typically, these methods are trained on finite, often biased datasets and may lack complete generalization in novel chemical spaces, resulting in inaccurate property predictions during the search process. This discrepancy can accumulate over iterations, leading to suboptimal molecular candidates or failure to meet the intended constraints. Moreover, the noise generated by these predictors may cause the search process to deviate from the optimal regions in the chemical or latent spaces, further reducing MO efficiency and effectiveness.
@@
To address the aforementioned MO issues, we explore the utilization of diffusion models in target-oriented MO. Diffusion models (Ho et al., 2020; Gong et al., 2024) have reaped remarkable success in other scientific disciplines, such as computer vision (Rombach et al., 2022), by generating high-quality data through a gradual denoising process that can capture complex distributions (Yang et al., 2023). Accordingly, we propose text-guided multi-property molecular optimization with a transformer-based diffusion language model (TransDLM). This is a novel approach that leverages a diffusion language model to iteratively yield word vectors of molecular SMILES strings and is guided by language descriptions. Due to SMILES strings’ deficiency in explicit semantic clarity of molecular structures and functional groups, we adopt standardized chemical nomenclature as informatively and intuitively molecular semantic representation. To differentiate generation from scratch, we sample molecular word vectors from the token embeddings of source molecules encoded by a pre-trained language model, which collectively and significantly retains the original molecules’ core scaffolds. Instead of relying on an external property predictor, we encode the molecular structure and property information using a pre-trained language model, implicitly embedding property requirements into textual descriptions, guiding diffusion and mitigating error propagation. This encoded representation serves as a guiding signal during the diffusion process. By embedding the desired molecular properties through the language model, the diffusion process is trained on the molecule’s structure and the specific properties we aim to optimize. Furthermore, we successfully optimized the binding selectivity of xanthine amine congener (XAC) from A2AR to A1R, which are both adenosine receptors, using TransDLM, indicating its broader applicability in optimizing ligand–receptor interactions in drug discovery. The TransDLM approach’s advantages are as follows:
@@
  • TransDLM outperforms state-of-the-art methods on the benchmark dataset in optimizing three ADMET [^1] Absorption, distribution, metabolism, excretion and toxicity properties, LogD, Solubility, and Clearance while maintaining the structural similarity over several metrics. And TransDLM successfully biases XAC’s selectivity preference from A2AR to A1R.
Figure 2. The framework of TransDLM. (a) Our optimization method disassembles the source molecules into fragments using the RECAP. Then, it generates textual descriptions utilizing IUPAC names. (b) The source and target SMILES strings are tokenized. They are then encoded into embeddings through a pre-trained language model, as well as the textual descriptions. (c) The embedded source and target molecules serve as initial and target representations during denoising in diffusion, respectively. The embeddings of textual descriptions guide the denoising process.
@@ 2.1. Molecular optimization
Guided search-based methods.   Guided search-based methods explore target molecules within the molecules’ chemical or latent space learned through encoder–decoder models using molecular property predictors or statistical models for guidance (Hoffman et al., 2022). For example, Jin et al. (Jin et al., 2018) introduced a junction tree variational autoencoder (JTVAE), which initially decomposes a molecular graph into a junction tree, and both the junction tree and molecular graph are mapped to latent embeddings. Finally, gradient ascent is used to identify embeddings with enhanced property scores within the latent space. Zhou et al. (Zhou et al., 2019) proposed the MolDQN model, which integrates reinforcement learning with chemical rules to guarantee molecular validity. The model frames molecule modification as a Markov Decision Process (MDP) and uses deep Q-networks (Hester et al., 2018) to address it. Nevertheless, these guided search-based methods rely on external property predictors to iteratively optimize molecular properties, inevitably introducing errors and noise into the optimization process.
@@
Molecular mapping-based methods.   Two molecules with moderate structural difference and significant chemical dissimilarities are considered as a matched molecular pair (MMP), where the one with inferior properties is regarded as the source molecule and the other with desired qualities is the target molecule. Molecular mapping-based method is usually executed using medicinal chemistry transformation rules derived from MMPs, applying the knowledge and expertise of medicinal chemists. For instance, He et al. (He et al., 2021) designed a model based on Seq2Seq architecture to handle property variations between MMPs. The model uses source molecules and property constraints as inputs and is trained to produce target molecules as outputs. Maragakis et al. (Maragakis et al., 2020) introduced DESMILES, which utilizes source molecular fingerprints as inputs to generate corresponding SMILES representations. Initially, the transformation from molecular fingerprints to SMILES is pre-trained on a large dataset, and the model is subsequently fine-tuned using the MMPs. Inspired by natural language processing (NLP), Fan et al. (Fan et al., 2022) developed a semi-supervised approach named BT4MolGen. This method involves training a pseudo-labeled data generation model using MMPs. Then, the generated pseudo-labeled data along with labeled data are employed to train a forward molecular optimization model, addressing the data sparsity issue.
@@ 2.2. Language generation with diffusion models
Diffusion models have been extremely successful in generating content in continuous domains, particularly in the images and audio fields (Yang et al., 2023). Stable Diffusion (Rombach et al., 2022) and AudioLDM (Liu et al., 2023) are categorized as diffusion models in continuous domains. They introduced random noises into latent variables based on latent diffusion models (LDMs) and reverse the process through a series of denoising steps to learn data generation. However, differences in data structure between languages and images exist, since languages are discrete while images exist in continuous domains. In order to tackle this problem, certain approaches retain the discrete nature of text and extend diffusion model to handle discrete data (Reid et al., 2022; He et al., 2023). On the contrary, other methods utilize embedding layers to map the text to a continuous representation domain, thereby preserving the continuous diffusion steps (Li et al., 2022; Gong et al., 2023; Yuan et al., 2022). Our research aligns with the latter strategy, focusing on word vectors and substantially expanding the functionality and feasibility of diffusion models in SMILES-based MO.
@@ 3.1. Overall framework
With the solid foundation paved by conditional language generation tasks, exemplified by DiffuSeq (Gong et al., 2023) and SeqDiffuSeq (Yuan et al., 2022), text-guided MO has become feasible. In this work, we aim to optimize a molecule based on a given textual description. Formally, let $C=[w_{0},w_{1},\dots,w_{m}]$ denote the input description, where $w_{i}$ is the $i$-th word and $m$ is the sequence length. The objective is to learn a model $f_{\theta}(\cdot)$ that takes $C$ and an initial molecule $M_{0}$ as input and outputs the optimized molecule: $M=f_{\theta}(M_{0},C)$. Overall, our TransDLM is composed of four pivotal processes: embedding, noising, denoising, and rounding. As shown in Fig. 2, these processes work in tandem to produce the desired optimized molecules.
Embedding. The text sequence $W=[w_{0},w_{1},\dots,w_{n}]$ is converted into embeddings via $\text{Emb}(W)=[\text{Emb}(w_{0}),\text{Emb}(w_{1}),\dots,\\ \text{Emb}(w_{n})]\in\mathbb{R}^{d\times n}$, where $d$ is the embedding dimension and $n$ is the maximum sequence length.
Noising. The initial latent representation $x_{0}$ is drawn from a Gaussian distribution centered at $\text{Emb}(W)$: $x_{0}\sim\mathcal{N}(\text{Emb}(W),\sigma_{0}\mathbf{I})$. Noise is then gradually added over $T$ diffusion steps until $x_{T}\sim\mathcal{N}(0,\mathbf{I})$. The transition from $x_{t-1}$ to $x_{t}$ is defined as follows:
@@
where $\beta_{t}$ controls the noise level at step $t\in(0,T]$.
Denoising. This process begins with an encoded source molecule, and sequentially samples $x_{t-1}$ from $x_{t}$, gradually reconstructing the desired content. A pre-trained language model maps $C$ to latent embedding $\mathbf{C}\in\mathbb{R}^{d_{1}\times m}$, where $d_{1}$ represents the embedding dimension of the language model output. Typically, a neural network is trained to predict $x_{t-1}$ given $x_{t}$. To improve the accuracy of denoising $x_{t}$ towards specific word vectors, a neural network $f_{\theta}(\cdot)$ is trained to predict $x_{0}$ directly from $x_{t}$, yielding the predicted target molecule embedding $\hat{x}_{0}=f_{\theta}(x_{t},t,\mathbf{C})$. Thus, the denoising transition from $x_{t}$ to $x_{t-1}$ can be expressed as:
(2) $p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t,\mathbf{C}),\textstyle\sum_{t}),$
(3) $\mu_{\theta}(x_{t},t,\mathbf{C})=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}f_{\theta}(x_{t},t,\mathbf{C})+\frac{\sqrt{\alpha_{t}}(1-\overline{\alpha}_{t-1})}{1-\overline{\alpha}_{t}}x_{t},$
where $\sum_{t}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t}$ is the covariance matrix, and $\overline{\alpha}_{t}=\prod^{t}_{s=0}(1-\beta_{s})$, facilitating the iterative sampling of $x_{t-1}$ from $x_{t}$ and yielding $\hat{x}_{0}$.
Rounding. The predicted latent embedding $\hat{x}_{0}$ is converted back to SMILES by matching each column of the vector to the nearest token embedding in L-2 distance. As a result, denoising and rounding together transform any initial distribution into a coherent SMILES string output.
In short, TransDLM integrates these four stages to iteratively transform a source molecule into a target molecule which meets the specified textual description.